Compute the horizontal range, on level ground, for a projectile with initial speed 35.0m/s and initial inclination 75.00 from the horizontal.


1
Expert's answer
2022-06-28T08:19:05-0400

The horizontal range is given by

"R=\frac{v_0^2\sin2\theta}{g}"

"R=\frac{35.0^2\sin(2*75.0^\circ)}{9.8}=62.5\:\rm m"